As we've seen, astronauts measure their mass by measuring the period of oscillation when sitting in a chair connected to a spring. The Body Mass Measurement Device on Skylab, a 1970s space station, had a spring constant of 606 N/m . The empty chair oscillated with a period of 0.802 s .

What is the mass of an astronaut who oscillates with a period of 2.28 s when sitting in the chair?
Answer #1

Time Period in SHM motion is given by:

T = 2*pi*sqrt (m/k)

When empty chair oscillates, then

m = m_c = mass of chair

k = spring constant = 606 N/m

T = time period of empty chair = 0.802 sec

In this case

0.802 = 2*pi*sqrt (m_c/k)

Now when astronaut sitting on chair oscillates, then

m = m_c + m_a = mass of chair + mass of astronaut

T = time period of astronaut in chair = 2.28 sec

So, in this case

2.28 = 2*pi*sqrt ((m_c + m_a)/k)

Now divide both equation

0.802/2.28 = sqrt (m_c/(m_c + m_a))

square both side

0.12373 = m_c/(m_c + m_a)

0.12373*m_a = m_c - 0.12373*m_c

m_c = [0.12373/(1 - 0.12373)]*m_a

m_c = 0.1412*m_a

Now Using this value in 2nd equation

2.28 = 2*pi*sqrt ((m_c + m_a)/k)

2.28 = 2*pi*sqrt ((0.1412*m_a + m_a)/k)

Using value of k and squaring both sides

[2.28/(2*pi)]^2 = 1.1412*m_a/606

m_a = 606*2.28^2/(1.1412*(2*pi)^2)

m_a = mass of astronaut = 69.92 kg

Please Upvote